Hoping someone will help me with a few more. I need to see the steps:)
Verify:
a)sinx(sinx+cscxcos^2x)=1
b)tanx((cotx/cosx)-sinx)=cosx
c)(1+cosx)(1-cosx)=sin^2x

Answer by stanbon(75887)

Verify:
a)sinx(sinx+cscxcos^2x)=1
---
sin^2(x) + (1)cos^2(x) = 1
---
sin^2(x) + cos^2(x) is always 1, so you get:
1 = 1
-----------------------
b)tanx((cotx/cosx)-sinx)=cosx
---
tan(x)[(1/tan(x))/cos(x) - sin(x)) = cos(x)
---
(1/cos(x) - sin(x) = cos(x)
Multiply thru by cos(x) to get:
1 - cos(x)sin(x) = cos^2(x)
This is not an identity.
The left side of the original eq. does not equal the right side.
-------------------------------------------------

c)(1+cosx)(1-cosx)=sin^2x
-----
1-cos^2(x) = sin^2(x)
sin^2(x) = sin^2(x)
=============================
Cheers,
Stan H.